# LLM2TEA: Agentic AI Designer Finds Innovative Objects with Generative Evolutionary Multitasking

### v2

@@ Abstract
In this paper, we introduce LLM-driven MultiTask Evolutionary Algorithm (LLM2TEA), the first agentic AI designer within a generative evolutionary multitasking (GEM) framework that promotes the crossover and synergy of designs from multiple domains, leading to innovative solutions that transcend individual disciplines. Of particular interest is the discovery of objects that are not only innovative but also conform to the physical specifications of the real world in science and engineering. LLM2TEA comprises a large language model to initialize a population of genotypes (defined by text prompts) describing the objects of interest, a text-to-3D generative model to produce phenotypes from these prompts, a classifier to interpret the semantic representations of the objects, and a physics simulation model to assess their physical properties. We propose several novel LLM-based multitask evolutionary operators to guide the search toward the discovery of high-performing practical objects. Experimental results in conceptual design optimization validate the effectiveness of LLM2TEA, revealing from 97% to 174% improvement in the diversity of innovative objects compared to the present text-to-3D generative model baseline. In addition, more than 73% of the generated designs have better physical performance than the top 1% percentile of the designs generated in the baseline. Moreover, LLM2TEA generates designs that are not only aesthetically creative but also functional in real-world applications. Several of these designs have been successfully 3D-printed, emphasizing the proposed approach’s capacity to transform AI-generated outputs into tangible physical objects. The designs produced by LLM2TEA meets practical requirements while showcasing creative and innovative features, underscoring its potential applications in complex design optimization and discovery.
@@
Agentic AI Designer, LLMs and text-to-3D Generative AI, Engineering Design Optimization and Discovery, LLM Evolutionary Optimizers, Evolutionary MultitaskingI Introduction
@@
Generative Artificial Intelligence (GenAI) has showcased remarkable capabilities in creating diverse creative objects from free-form natural language prompts [1]. Examples include the synthesis of creative texts, images, audio, and videos through tools such as SARD [2], Stable Diffusion [3], Make-An-Audio [4], and Veo 2.0 [5]. Over the past few years, a surge of research in GenAI has focused on enabling the production of high-quality digital media that closely aligns with user preferences through tailored prompts. However, despite these widely publicized advancements in digital media, there remains a notable gap in the application of GenAI for driving progress in complex engineering design, with comparatively fewer reports addressing its potential in the science and engineering domains.
@@
In the realm of traditional engineering design, which is the primary focus of our work, we have shown the promising potential of GenAI in conceptual aerodynamic design [6, 7]. The major challenge in applying GenAI to science and engineering that we have identified lies not only in generating digital objects that exhibit high novelty but also in ensuring that these objects can be translated into tangible physical solutions that adhere to natural principles [8] or meet specific problem contexts and design specifications [9]. Our early research studies have attempted to address this challenge by advancing GenAI for engineering design optimization [10, 11, 12], where preliminary results reveal encouraging progress, particularly with the use of text prompt optimization to enhance aerodynamic performance [6] while ensuring semantic conformity [7]. These efforts underscore the potential of GenAI to bridge the gap between digital innovation and practical applicability, particularly in engineering disciplines.
@@
To this end, we introduce LLM-driven MultiTask Evolutionary Algorithm (LLM2TEA), the first agentic AI designer within a generative evolutionary multitasking (GEM) framework designed to facilitate the cross-pollination of designs across multiple domains. LLM2TEA aims to uncover innovative solutions that transcend disciplinary boundaries while adhering to natural principles and meeting design specifications. LLM2TEA comprises an LLM-driven multitask evolutionary algorithm, a text-to-3D generative model for phenotype production, a classifier for interpreting semantic representations, and a physics simulation model to assess the physical properties of the generated objects. By solving multiple tasks simultaneously, where each task focuses on a single-domain search, LLM2TEA improves design diversity by exploring beyond individual domains and into the intersections of domains to search for cross-domain designs with cross-domain knowledge transfer.
@@
We propose several novel LLM-based multitask evolutionary operators, including new instruction sets to guide the evolutionary processes. This includes an initialization operator for generating initial populations of genotypes and multitask mating operators like crossover, mutation, and selection to enable an efficient exchange of traits across domains as evolution progresses with the generation of new offspring. The population of genotypes serves as guides for the text-to-3D generative model for phenotype synthesis. The survival of the best-performing designs is assessed based on the practicality of these phenotypes with criteria including physical properties and semantic meanings. To summarize, the contributions of this paper are as follows:
@@
This paper introduces LLM2TEA, the first agentic AI designer within a generative evolutionary multitasking framework that promotes the cross-pollination of designs from multiple domains, leading to innovative solutions that transcend individual disciplines.
@@
An LLM-driven multitask evolutionary algorithm with operators synthesizing creative text prompts as guides for an auxiliary text-to-3D generative model to produce innovative 3D objects. LLM2TEA is the first GEM method to leverage LLM capabilities for complex design optimization and discovery of digital objects that exhibit high novelty while translatable into tangible physical solutions that adhere to natural principles and meet specific problem contexts and design specifications.
@@
A series of complex design and discovery tasks mimicking real-world design optimization scenarios are demonstrated to verify the efficacy of the proposed LLM2TEA. Our experimental results show that LLM2TEA excels in uncovering innovative designs and/or improved aerodynamic performance compared to the present text-to-3D generative model as a baseline. We evaluate with a proposed innovation score metric and show that our proposed approaches led to more than of innovative designs found above the mean baseline.
We demonstrate that some of the designs discovered with LLM2TEA can be 3D printed using off-the-shelf stereolithography (SLA) printing technology. As previously highlighted, while a direct use of the text-to-3D generative model tends to generate designs that are infeasible in practice, the findings showed that LLM2TEA, on the other hand, uncovers designs that are compatible with SLA printing, suggesting it as a notable approach for practical rapid design prototyping.
@@
The remainder of the paper is organized as follows: Section II reviews related work from the literature and provides an overview of key background concepts. Section III introduces LLM2TEA in detail. Section IV presents the experimental setup, and the results and discussions are detailed in Section V. Finally, Section VI concludes this paper and highlights the future direction.
@@ II Related Works
In this section, we provide an overview of related work. First, we discuss the applications of text-to-X generative models (where X represents various outputs such as languages, images, or 3D shapes) in creative and engineering conceptual design domains. Next, we review prompt engineering and optimization techniques, primarily used to generate suitable prompts for text-to-X generative models. Finally, we present the recent advancements in the evolutionary multitasking (EMT) paradigm and discuss the possibility of handling cross-domain optimization tasks simultaneously.
@@ II-A Text-to-X Generative Models in Creative & Engineering Conceptual Design
II-B Prompt Engineering and Optimization
We observe that augmenting prompts impacts the feasibility of generated artifacts in text-to-X generative models. Consequently, designing effective prompts is crucial for leveraging the full potential of generative models to solve various tasks. Several works proposed prompting techniques that introduced handcrafted prompt templates to accomplish target tasks better [16, 17]. On the other hand, prompt optimization techniques introduced optimizers to search for the optimal prompt by utilizing optimizers [18, 19, 20].
In recent years, prompt optimization techniques have also been applied in engineering design optimization [6]. In that work, the authors proposed a fully automated evolutionary design optimization framework incorporating evolutionary strategy and performance-based guidance to seek novel 3D car designs in Shap-E text-to-3D generative model. However, the design search in the text-prompt space can sometimes lead to many ill-defined designs as Shap-E was not trained specifically on a dataset of car shapes. To mitigate such issues, vision-language models [21, 22] were introduced to the framework that penalizes ill-defined designs [7]. The research demonstrated the possibility of optimizing generated designs via evolutionary methods that optimize only the prompts to satisfy both practicality and visual requirements. However, the optimized design often still highly resembles car shapes as the optimization objective solely focuses on seeking car designs.
Due to the advancements in language comprehension and generation capabilities, LLMs are also considered effective tools for generating prompts and can even be employed as optimizers. For instance, Yang et al. [23] directly utilized an LLM as the optimizer for the prompt optimization task, demonstrating the feasibility and effectiveness of this approach. Guo et al. [24] combined evolutionary algorithms with an LLM to enhance prompt optimization, leveraging the strengths of both methods. Additionally, Yang et al. [25] employed evolutionary multi-objective optimization to search for Pareto-optimal prompts. However, most of the aforementioned methods focus on a single task. As discussed in Section I, we posit that incorporating cross-domain learning into the optimization process can lead to prompts that may yield innovative design solutions. Therefore, we propose to achieve prompt optimization by employing evolutionary multitasking techniques. This approach leverages knowledge from multiple domains to generate prompts that are not only effective but also capable of discovering novel designs.
@@ II-C Evolutionary Multitasking
Traditionally, the primary focus of conventional evolutionary algorithms has been on generating a set of solutions for a single optimization task [26]. However, inspired by the human brain’s capability for parallel task processing, the EMT paradigm has recently emerged as a population-based search methodology specifically designed to address multiple optimization tasks concurrently[27]. In contrast to traditional evolutionary algorithms, EMT exploits latent synergies between distinct yet related optimization tasks, thereby enhancing search performance in terms of solution quality. Over the past few years, several successful applications of EMT have been documented across various domains, including manufacturing process design [28], mobile edge computing [29], and vehicle routing [30, 31]. These examples highlight EMT’s potential in efficiently solving complex, multi-faceted problems by leveraging cross-task knowledge sharing.
There has been a growing interest in LLM as evolutionary optimizers to tackle tasks encompassing both traditional optimization tasks [32, 33] and prompt optimization tasks [24]. Despite these advancements, there has been no documented work on the LLM performing EMT operations for optimization purposes. In this paper, we demonstrate that LLMs are compatible with EMT to not only facilitate the optimization process but also promote the discovery of innovative and creative solutions.
III Large Language Model-Driven MultiTasking Evolutionary Algorithm
In this section, we provide a detailed description of the proposed LLM2TEA. First, we present the multitask optimization problem of interest. Next, we delve into each distinctive feature of the LLM2TEA, describing its role in the multitask evolutionary algorithm.
@@ III-A Preliminaries
Let the denote the search space of sequential tokens, where is the maximum sequence tokens length and is the vocabulary size. Given a set of tasks , the LLM constructs a genotype (prompt) () relevant to the -th task using the following template:
@@
Here, user provides a target domain “” that defines the task of interest (e.g., “” could be set to “car” or “airplane”). The variable represents the part of the prompt that LLM2TEA will optimize. Based on this setting, the single-objective function for the -th task is formulated as follows:
@@
where and control the exploration of designs within and near domain boundary, and solutions trade-offs among the multiple criteria are of interest. The setting of these objective functions ensures that the optimized prompt generates shapes that not only meet the target physical criteria but also resemble the conceptual elements of the target domain represented by .
@@ III-B LLM2TEA Overview
The LLM2TEA presented in Fig. 1 comprises an LLM that constructs prompts from a set of instructions (depicted in Fig. 2) containing the user specifications embedded in . The generated prompts are served to a text-to-3D generative model that synthesizes designs, which are evaluated against a set of criteria for computing the fitness of the designs with respect to the target objective. Details of the key algorithmic features of LLM2TEA for different proposed approaches in complex design and discovery are discussed in this section.
@@ III-B1 Population Initialization
LLM2TEA begins with instructing an LLM to generate an initial population of genotypes, each genotype being a text prompt based on free-form natural language as discussed in Section III-A. Specifically, the LLM is presented with a description of the optimization problem (shown in Fig. 2(a)) and is instructed to generate a population of genotypes (shown in Fig. 2(b)). These genotypes are served to a text-to-3D generative model for synthesizing the corresponding 3D designs.
@@ III-B2 Offspring Generation
It has been demonstrated that the LLM’s in-context learning capability integrates historical objective evaluation data, enhancing guidance [32]. Leveraging on this finding, our proposed LLM2TEA provides the LLM with the current population’s genotypes and instructs it to reflect on them (as shown in Fig. 2(c)) before generating the offspring. This reflection ensures context-aware execution of the assigned mating operations, aligning genetic variation with the algorithm’s objectives.
@@
During offspring generation, the LLM is instructed to select two parent genotypes from the current population and execute crossover and mutation operations. Conventional EAs typically employ simple genetic crossover operators [34] or probabilistic models, such as Gaussian distributions [35], as mutation operators [36]. However, these traditional crossover and mutation techniques exhibit two main limitations that could inhibit the effectiveness of prompt optimization techniques. Firstly, meaningful prompts in natural language often entail implicit constraints inherent in language sequences. Traditional crossover and mutation operators fail to account for these implicit constraints, potentially resulting in nonsensical prompts that may not effectively generate reasonable outputs. Secondly, considering that an LLM contains extensive prior knowledge, leveraging it for crossover and mutation [37] can facilitate the generation of more coherent and contextually appropriate prompts by utilizing this embedded knowledge.
@@
Building on the above considerations, LLM plays the role of an evolutionary operator to facilitate the mating process. This process involves selecting parents, followed by crossover and/or mutation operations to construct a genotype, and finally, the generation of a design conditioned by the constructed genotype. In LLM2TEA, we propose self-mating, same-domain mating, and cross-domain mating operators, which are described as follows:
@@
Self-Mating: The LLM is instructed to select a parent genotype from the current population and construct a new genotype based on the context of the selected parent.
@@
Same-Domain Mating: An instruction is provided to the LLM to randomly select two parent genotypes belonging to the same domain to instantiate the genotype template by picking some words from both parents. The model mutates this instantiated genotype by randomly replacing some words with other words in the genotype.
@@
Cross-Domain Mating: Instructed to randomly select two parent genotypes associated with two different domains, the LLM will instantiate the genotype template by picking some words from both parents. The model then mutates the instantiated genotype by randomly replacing some words with other words.
@@
During the mating process, if both selected parents belong to the same domain, the same-domain mating operator (shown in Fig. 2 (e)) kicks in; otherwise, a cross-domain mating operator (shown in Fig. 2 (f)) takes effect. Additionally, the self-mating operator (shown in Fig. 2 (d)) involving a single parent is triggered with a certain probability. The resulting genotypes are served to a text-to-3D generative model, thus synthesizing new 3D designs. The generated 3D design and its associated prompt are the objects of an offspring genotype. All the offspring genotypes will form the offspring population and are evaluated against the physical, geometric, and visual criteria. Subsequently, the offspring population is combined with the current population as the joint population. Any duplicate genotypes found in this joint population will be removed in the current generation population, retaining the unique genotype originating from the offspring population. The resulting joint population undergoes selection pressure based on the surviving genotypes.

### v3

# LLM2TEA: An Agentic AI Designer for Discovery with Generative Evolutionary Multitasking

@@ Abstract
This paper presents LLM2TEA, a Large Language Model (LLM) driven MultiTask Evolutionary Algorithm, representing the first agentic AI designer of its kind operating with generative evolutionary multitasking (GEM). LLM2TEA enables the crossbreeding of solutions from multiple domains, fostering novel solutions that transcend disciplinary boundaries. Of particular interest is the ability to discover designs that are both novel and conforming to real-world physical specifications. LLM2TEA comprises an LLM to generate genotype samples from text prompts describing target objects, a text-to-3D generative model to produce corresponding phenotypes, a classifier to interpret its semantic representations, and a computational simulator to assess its physical properties. Novel LLM-based multitask evolutionary operators are introduced to guide the search towards high-performing, practically viable designs. Experimental results in conceptual design optimization validate the effectiveness of LLM2TEA, showing to improvements in the diversity of novel designs over the current text-to-3D baseline. Moreover, over of the generated designs outperform the top of designs produced by the text-to-3D baseline in terms of physical performance. The designs produced by LLM2TEA are not only aesthetically creative but also functional in real-world contexts. Several of these designs have been successfully 3D printed, demonstrating the ability of our approach to transform AI-generated outputs into tangible, physical designs. These designs underscore the potential of LLM2TEA as a powerful tool for complex design optimization and discovery, capable of producing novel and physically viable designs.
@@
Agentic AI Designer, Large Language Models, text-to-3D Generative AI, Generative Evolutionary Multitasking, Engineering Design Optimization and Discovery, LLM Evolutionary Optimizers, Evolutionary MultitaskingI Introduction
@@
Generative Artificial Intelligence (GenAI) has demonstrated remarkable capabilities in producing diverse creative designs from free-form natural language prompts [1]. Notable examples include the synthesis of creative texts, images, audio, and videos through tools such as SARD [2], Stable Diffusion [3], Make-An-Audio [4], and Veo 2.0 [5]. Over the past few years, a surge of research in GenAI has focused on enabling the production of high-quality digital media that closely aligns with user preferences through the use of tailored prompts. Despite these widely publicized advancements in digital media, a notable gap remains in applying GenAI to complex engineering design, with comparatively fewer reports addressing its potential within the science and engineering domains.
@@
In the realm of traditional engineering design, which is the primary focus of this work, the promising potential of GenAI in conceptual aerodynamic design has been demonstrated [6, 7]. The major challenge identified in applying GenAI to science and engineering lies not only in generating digital designs exhibiting high novelties, but also in ensuring that the designs can translate into tangible physical solutions that adhere to natural principles [8] or satisfy specific problem contexts and design specifications [9]. Our early research studies have attempted to address this challenge by advancing GenAI for engineering design optimization [10, 11, 12]. Preliminary results leveraging conventional optimization strategies indicate encouraging progress, particularly through the use of text prompt search to improve aerodynamic performance [6] while maintaining semantic conformity [7]. These efforts underscore the potential of GenAI to bridge the gap between digital innovation and practical applicability, particularly in engineering disciplines. However, navigating the very high-dimensional language search space often demands an integration of symbolic, statistical, and experiential knowledge to support broader generalization and knowledge transfer. This imperative highlights the need for agentic AI approaches that can engage in objective-directed reasoning and guidance as well as adaptive problem-solving.
@@
To this end, a large language model (LLM) driven MultiTask Evolutionary Algorithm (LLM2TEA), is introduced as the agentic AI designer implemented with generative evolutionary multitasking (GEM), structured to facilitate the cross-pollination of solutions across multiple domains. The objective of LLM2TEA is the discovery of novel solutions that transcend disciplinary boundaries while adhering to natural principles and meeting user specifications. LLM2TEA is composed of an LLM-guided multitask evolutionary algorithm, a text-to-3D generative model responsible for phenotype production, a classifier for interpreting semantic representations, and a physics simulation model to evaluate the physical properties of the generated designs. By solving multiple tasks simultaneously, where each task corresponds to a single-domain search, the exploration of design space is expanded beyond isolated domains and into their intersections, allowing for the emergence of cross-domain designs enabled by knowledge transfer.
@@
Several novel LLM-based multitask evolutionary operators are introduced, incorporating new instruction sets to guide the evolutionary processes. This includes an initialization operator for generating an initial population of genotypes and multitask mating operators, such as crossover, mutation, and selection, to enable the efficient exchange of traits across domains as evolution progresses with the generation of offspring. The resulting genotype population serves as input to the text-to-3D generative model for synthesizing phenotypes. The survival of the best-performing designs is evaluated based on the practical viability of these phenotypes, considering criteria that include physical properties and semantic meanings. In summary, the contributions of this paper are as follows:
@@
This paper introduces LLM2TEA, a novel agentic AI designer with a generative evolutionary multitasking (GEM) that promotes the crossbreeding of designs from multiple domains, leading to novel solutions that transcend individual domains.
@@
LLM2TEA is an LLM-driven multitask evolutionary algorithm with operators synthesizing creative text prompts as guides to an auxiliary text-to-3D generative model in producing corresponding novel 3D designs. LLM2TEA represents the first method operating with the GEM paradigm to leverage the capabilities of LLMs for complex design optimization and the discovery of novel digital designs. These digital designs not only exhibit high novelty but are also translatable into tangible physical solutions that conform to natural laws and fulfill specific problem contexts and design specifications.
@@
A series of complex design and discovery tasks, modeled after real-world design optimization scenarios, is conducted to assess the efficacy of the proposed LLM2TEA. Experimental results indicate that LLM2TEA consistently outperforms the baseline text-to-3D generative model in uncovering novel designs and/or enhancing aerodynamic performance. The proposed novelty score metric is employed for evaluation, with findings showing that over of generated designs are identified as novel, exceeding the mean performance of the baseline.
It is demonstrated that several designs discovered through LLM2TEA are physically realizable and successfully 3D printed using off-the-shelf stereolithography (SLA) technology. As previously noted, direct outputs from baseline text-to-3D generative models often result in designs that are highly fragmented for physical fabrication. In contrast, the findings indicate that LLM2TEA is capable of identifying physically feasible designs suitable for SLA printing, highlighting its potential as a viable approach for practical rapid design prototyping.
@@
The remainder of the paper is organized as follows: Section II reviews related work from the literature and provides an overview of key background concepts. Section III introduces LLM2TEA in detail. Section IV presents the experimental setup, and the results and discussions are detailed in Section V. Finally, Section VI concludes this paper, highlighting some potential future directions.
@@ II Related Works
This section presents an overview of related work. First, the applications of text-to-X generative models (where X represents various outputs such as languages, images, or 3D shapes) are reviewed in the context of creative and engineering conceptual design. Next, prompt engineering and search techniques are examined, with a focus on methods to generate effective prompts for text-to-X generative models. Finally, recent advancements in the evolutionary multitasking (EMT) paradigm are discussed, with an emphasis on its potential to address cross-domain optimization tasks simultaneously.
@@ II-A Text-to-X Generative Models in Creative & Engineering Conceptual Design
II-B Prompt Engineering and Search
It has been observed that prompt augmentation significantly influences the feasibility and quality of artifacts generated by text-to-X generative models. Consequently, designing effective prompts is crucial for fully harnessing the full potential of these generative models for solving diverse tasks. Several studies have proposed prompting techniques based on handcrafted prompt templates to better align the outputs of the model with target objectives [16, 17]. In contrast, prompt search methods have considered optimization approaches to finding optimal prompts through iterative refinement [18, 19, 20].
In recent years, prompt search techniques have also been applied in engineering design optimization [6]. Rios et al. proposed a fully automated evolutionary design optimization framework incorporating evolutionary strategies with performance-based guidance to seek for novel 3D car designs using the Shap-E text-to-3D generative model. However, searching within the text-prompt space can often lead to ill-defined design outputs, as Shap-E may not be explicitly trained on a dataset focused on car shapes. To address this limitation, vision-language models [21, 22] were introduced into the framework to penalize semantically incoherent or structurally implausible designs [7]. The research demonstrated the feasibility of searching effective text prompts to generate designs that meet both practical constraints and visual requirements with evolutionary approaches. However, the resulting design often retained strong resemblances to conventional car shapes, since the optimization process was narrowly targeted at car design objectives.
Due to recent advancements in language comprehension and generation capabilities, LLMs have become effective tools not only for generating prompts but also for serving as optimizers. For instance, Yang et al. [23] directly utilized an LLM as the optimizer in the prompt optimization task, demonstrating both the feasibility and effectiveness of such an approach. Guo et al. [24] combined evolutionary algorithms with an LLM to enhance prompt optimization, leveraging the strengths of both methods. However, most existing approaches primarily focus on optimizing prompts for a single task. As discussed in Section I, it is believed that incorporating cross-domain learning into the design optimization process can identify prompts that lead to novel design solutions. Therefore, prompt search via evolutionary multitasking techniques is proposed. Such an approach leverages knowledge from multiple domains to explore a broader design space that extends beyond individual domains and embraces its overlap, prompting the discovery of novel cross-domain designs through knowledge transfer.
@@ II-C Evolutionary Multitasking
Recently, the evolutionary multitasking (EMT) paradigm has emerged as a population-based search methodology designed to handle multiple optimization tasks concurrently [25]. In contrast to traditional evolutionary algorithms, EMT exploits latent synergies between distinct yet related optimization tasks, thereby enhancing convergence performance. Over the past few years, several successful applications of EMT have been documented across various domains, including manufacturing process design [26], mobile edge computing [27], and vehicle routing [28, 29]. These examples highlight the potential of EMTs for efficiently solving complex, multifaceted problems by leveraging cross-task knowledge sharing.
While there is surging research in using LLMs as evolutionary optimizers to address both traditional optimization [30, 31] and prompt optimization tasks [24], no documented studies have explored exploiting the contrasting characteristics of multiple domains to enable crossbreeding of solutions with evolutionary multitasking, resulting in the discovery of novel artifacts with hybrid features. This paper aims to investigate this possibility, proposing a novel prompt search technique that operates with generative evolutionary multitasking (GEM). The proposed method not only demonstrates effective optimization but also promotes the discovery of novel solutions.
III Agentic AI Designer With Generative Evolutionary Multitasking (GEM)
In this section, the proposed LLM2TEA is presented. First, the multitask optimization problem of interest is described. Next, each distinctive feature of the LLM2TEA, describing its role in the multitask evolutionary algorithm, is examined in detail.
@@ III-A Preliminaries
Let denote the search space of sequential tokens, where is the maximum sequence tokens length and is the vocabulary size. Given a set of tasks , the LLM constructs a genotype (prompt) () relevant to the -th task using the following template:
@@
Here, a user provides a target domain “” that defines the task of interest (e.g., “” as “car”, “airplane” or others). The variable represents the part of the prompt that LLM2TEA will optimize. Based on this setting, the single-objective function for the -th task is formulated as follows:
@@
where and control the exploration of designs within and near the domain boundary, and solutions trade-offs among the multiple criteria are of interest. The setting of these objective functions ensures that the optimized prompt generates shapes that not only meet the target physical criteria but also resemble the conceptual elements of the target domain represented by .
@@ III-B LLM2TEA Overview
Given the very high-dimensional nature of the language search space, assume the existence of a function that can construct the prompt , where denotes the historical context, is the maximum sequence context length and is the context vocabulary size. Further, the hidden Markov property is assumed to hold, which states that all necessary information to derive an outcome exists in the current state (or context in our case), thereby reducing . Under these assumptions, the capabilities of an LLM are utilized as the function to construct effective prompts that guide the text-to-3D model in generating aerodynamically performing vehicle designs. With these considerations, LLM2TEA (Fig. 1) comprises an LLM that constructs prompts from a set of instructions (depicted in Fig. 2) containing the user specifications embedded in . The generated prompts are passed to a text-to-3D generative model that synthesizes designs, which are then evaluated against a set of criteria to compute the fitness with respect to the target objective. Details of the key algorithmic features of LLM2TEA for various proposed approaches in complex design and discovery are elaborated upon in this section.
@@ III-B1 Population Initialization
LLM2TEA begins with instructing an LLM to generate an initial population of genotypes, each genotype being a text prompt based on free-form natural language as discussed in Section III-A. Specifically, the LLM is presented with a description of the optimization problem (shown in Fig. 2(a)) and is instructed to generate a population of genotypes (shown in Fig. 2(b)). These genotypes are then passed to a text-to-3D generative model for synthesizing the corresponding 3D designs.
@@ III-B2 Offspring Generation
It has been demonstrated that the in-context learning capability of LLM can incorporate historical objective evaluation data to enhance optimization guidance [30]. Building on this insight, the proposed LLM2TEA provides the LLM with the current genotype population and instructs it to reflect on them (as shown in Fig. 2(c)) prior to generating offspring. This reflection ensures the context-aware execution of the assigned mating operations, aligning genetic variation effectively with the objectives of the algorithm.
@@
During offspring generation, the LLM is instructed to select parent genotypes from the current population and undergo crossover and/or mutation operations. Conventional EAs typically employ simple genetic crossover operators [32] or probabilistic models, such as Gaussian distributions [33], as mutation operators [34]. However, these traditional crossover and mutation techniques exhibit two main limitations that could inhibit the effectiveness of prompt search techniques. Firstly, meaningful prompts in natural language often entail implicit constraints inherent in language sequences. Traditional crossover and mutation operators fail to account for these implicit constraints, potentially resulting in nonsensical prompts that may not effectively generate reasonable outputs. Secondly, considering that an LLM contains extensive prior knowledge, harnessing for crossover and mutation [35] can facilitate the generation of more coherent and contextually appropriate prompts by utilizing this embedded knowledge.
@@
Building on the above considerations, LLM plays the role of an evolutionary operator to facilitate the mating process. This process involves selecting parents, followed by crossover and/or mutation operations to construct a genotype, and finally, generating a design conditioned by the constructed genotype. In LLM2TEA, several LLM-based self-mating, same-domain mating, and cross-domain mating operators are proposed, which are described as follows:
@@
Self-Mating: Instruction is provided to the LLM to select a parent genotype from the current population and construct a new genotype based on its context.
@@
Same-Domain Mating: Instruction is provided to the LLM to randomly select two or more same-domain parent genotypes, constructing a new genotype by combining words from both. The LLM can mutate the new same-domain genotype further by randomly replacing some words.
@@
Cross-Domain Mating: Instruction is provided to the LLM to randomly select two or more parent genotypes from different domains by combining words from both. The LLM can mutate the new cross-domain genotype further by randomly replacing some words.
@@
During the mating process, if both selected parents belong to the same domain, the same-domain mating operator (shown in Fig. 2 (e)) kicks in; otherwise, a cross-domain mating operator (shown in Fig. 2 (f)) takes effect. Additionally, the self-mating operator (shown in Fig. 2 (d)) involving a single parent is triggered with a certain probability. The resulting genotypes are passed into a text-to-3D generative model, thereby synthesizing new 3D designs. The generated 3D design and its associated prompt are the designs of an offspring genotype. All the offspring genotypes will form the offspring population and are evaluated based on the physical and visual criteria. Subsequently, the offspring population is combined with the current population as the joint population. Any duplicate genotypes found in this joint population will be removed in the current generation population, retaining the unique genotype originating from the offspring population. The resulting joint population undergoes selection pressure based on the surviving genotypes.